Project

General

Profile

11132013

Nightly Observing Plan for MJD 56610 (starting evening of Nov 13, 2013)

Observers

Observer 1 Observer 2 Run Manager SISPI Expert on Call
Matt Wiesner David James / Marisa March Marisa March
ext. 515 ext. 516 or 630 338 9267

CTIO staff (Telescope Operators)

Day Support Night Assistant
Hernan Tirado Leonardo Paredes

DECam Configuration and Setup

SISPI Configuration File DES.ini
DECam constants DECAM:CURRENT
Proposal ID 2012B-0001
Image Processing (Calibration) NCSA
Image Processing (First Cut) NCSA

Almanac
Current local time 15:34 CLST     UT 18:34     SidTim 17:23
Sunset     20:23 CLST  23:23 UT   SidTim  333 deg  
Eve. -10   21:01 CLST  00:01 UT   SidTim  343 deg  Zeropoint; standards
Eve. -14   21:23 CLST  00:23 UT   SidTim  348 deg  Begin obstac
Eve. -18   21:45 CLST  00:45 UT   SidTim  354 deg  Last chance standards
Midpoint   01:27 CLST  04:27 UT   SidTim   49 deg  
Moonset    04:36 CLST  07:36 UT   SidTim   96 deg  
Morn -18   05:09 CLST  08:09 UT   SidTim  105 deg  Begin astronomical twilight
Morn Std   05:31 CLST  08:31 UT   SidTim  110 deg  End obstac; begin standards
Morn -10   05:53 CLST  08:53 UT   SidTim  116 deg  
Sunrise    06:31 CLST  09:31 UT   SidTim  125 deg  
Moon position (deg): RA 9.1, Dec 7.2 Phase: 0.768

Weather Conditions

*Tololo Weather Links .

http://www.meteochile.gob.cl/cna/irsat_alta.png

Clear

Observing Plan and Goals for the Night

  1. Afternoon calibrations script to use: prenight_calibration.json
  2. Start of night standards
  3. ObsTac
  4. End of night standards
  5. End of night calibrations

Start of Night Standards
UT Date = '2013/11/14 00:01' Local time = 21:01
SDSSJ0200-0000 ra 02:00:00 dec +00:00:00 secz 1.7 script to use: DES_std_sdssj0200_Yzirg.json
SDSSJ2300-0000 ra 23:00:00 dec +00:00:00 secz 1.2 script to use: DES_std_sdssj2300_Yzirg.json

End of Night Standards
UT Date = '2013/11/14 08:31' Local time = 05:31
SDSSJ1048-0000 ra 10:48:00 dec +00:00:00 secz 1.9 script to use: DES_std_sdssj1048_grizY.json
SDSSJ0958-0010 ra 09:58:00 dec -00:10:00 secz 1.5 script to use: DES_std_sdssj0958_grizY.json
SDSSJ0843-0000 ra 08:43:00 dec +00:00:00 secz 1.2 script to use: DES_std_sdssj0834_grizY.json

Notes from the 4PM Meeting and Additional Comments:

Run Manager's Meeting
-------------------------------
Hernan, Michael, Margaret, Marisa, Brian, Eric, Tom, Manual, others
Matt

Telescope and Camera
--------------------
Camera and Telescope OK. SISPI restarted. Trane people working on Chiller #1.
They are looking for a leak. The telescope is on the white spot.

Last Night's Data
-----------------
As expected Obstac started with shallow SN fields because of deadman switch due to
calling some recent data bad. Then wide-fields. z-band because of the Moon.
Generally good seeing.

Problem once in the afternoon and then late at night with the DECal computer.
SISPI interlock for DECal came up as broken. Had to reset the DECal computer.
Yesterday there were problems with network in the control room and with VPN.
Marisa and TELOPs suspect the DECal problem is network related.
This cost 5 or 10 minutes.
Also a shutter alarm.

Quick Reduce keeps failing on some sort of key error. Angelo has been contacted
but because of the problem with the VPN Angelo couldn't connect to CTIO.
DESDM
-----
Margaret is about to send out DQ for 11/9 and 10th. The 11th is being
processed. And she has all the data from yesterday.

Tonight
-------
Weather is clear.

Obstac is predicted to do tiling 4 in z -30 < RA < 0, but the Moon is really
bright, so to get it the seeing is going to have to be really good. We didn't
get it last night because of the SN shallow field deadman.
We might do X3 r-band tonight on deadman.